OpenStudy (jazzyfa30):

Match the function with its value. f(x) = x2 - 3x + 3 g(x) = 2x2 -1 g(w3)

OpenStudy (jazzyfa30):

@hartnn

hartnn (hartnn):

g (w3) or g (f(3)) ?? whats w3 there ?

OpenStudy (jazzyfa30):

function thingy w^3

hartnn (hartnn):

well, if its g(w^3) then you just replace x by w^3 in g(x) g (w^3) = 2 (w^3)^2 -1 = ...?

OpenStudy (jdoe0001):

$$ \large { f(x) = x^2 - 3x + 3\\ g(x) = 2x^2 -1\\ g(whatever) = 2(whatever)^2 -1 } $$

OpenStudy (jazzyfa30):

2 (w^3)^2 -1 = 2w^6-1????

hartnn (hartnn):

that would be correct .

OpenStudy (jazzyfa30):

f(4)

hartnn (hartnn):

to get f(4), just plug in x = 4 in f(x) f(4) = 4^2-3*4+3

OpenStudy (jazzyfa30):

17

hartnn (hartnn):

4^2 =16 16 -12 = 4 4+7 = 7 just 7

hartnn (hartnn):

**4+3 =7

OpenStudy (jazzyfa30):

ok Determine if the function's degree is even or odd and the number of real zeros. Select one: a. odd with 3 real zeros b. odd with no real zeros c. even with 4 real zeros d. even with 2 real zeros

hartnn (hartnn):

number of zeros = number of times the graph crosses x axis =... ?

OpenStudy (jazzyfa30):

1

hartnn (hartnn):

sure ? 1 in the positive side, plus one in the negative side = ... ?

OpenStudy (jazzyfa30):

0

hartnn (hartnn):

no, it crosses x-axis twice, so d) even with 2 real zeros...

OpenStudy (jazzyfa30):

oh ok

hartnn (hartnn):

new question, new post please :)

OpenStudy (jazzyfa30):

ok